### v5

# Designing Large Foundation Models for Efficient Training and Inference: A Survey

Abstract
This paper focuses on modern efficient training and inference technologies on foundation models and illustrates them from two perspectives: model and system design. Model and System Design optimize LLM training and inference from different aspects to save computational resources, making LLMs more efficient, affordable, and more accessible. The paper list repository is available at https://github.com/NoakLiu/Efficient-Foundation-Models-Survey.
{forest}
forked edges, for tree= grow=east, reversed=true, anchor=base west, parent anchor=east, child anchor=west, base=left, font=, rectangle, draw=hidden-draw, rounded corners, align=left, minimum width=4em, edge+=darkgray, line width=1pt, s sep=3pt, inner xsep=2pt, inner ysep=3pt, line width=0.8pt, ver/.style=rotate=90, child anchor=north, parent anchor=south, anchor=center, where level=1fill=quant-color!30, font=, where level=2fill=quant-color!30, font=, where level=3fill=quant-color!30, font=, where level=4fill=quant-color!30, font= [ Efficient
Foundation
Models [ Model Design [ Quantization [ Bitwise-Based [ 1-bit/2-bit CODE(0x5649a7648030)eaf ] [ 4-bit CODE(0x5649a7648030)eaf ] [ 8-bit CODE(0x5649a7648030)eaf ] ] [ Method-Based [ PTQ (Post-Training Quantization) CODE(0x5649a7648030)eaf ] [ QAT (Quantization-Aware Training) CODE(0x5649a7648030)eaf ] ] ] [ Distillation [ Soft Knowledge Distillation [ Classical KD CODE(0x5649a7648030)eaf ] [ Temperature-Scaled KD CODE(0x5649a7648030)eaf ] ] [ Hard Knowledge Distillation [ Hard Label KD CODE(0x5649a7648030)eaf ] [ Program-Aided Distillation (PaD) CODE(0x5649a7648030)eaf ] ] ] [ Pruning [ Unstructured Pruning CODE(0x5649a7648030)eaf ] [ Structured Pruning CODE(0x5649a7648030)eaf ] ] ] [ System Design [ KV Cache Compression CODE(0x5649a7648030)eaf ] [ Parallelism CODE(0x5649a7648030)eaf ] [ Contextual Sparsity CODE(0x5649a7648030)eaf ] [ Memory Management CODE(0x5649a7648030)eaf ] [ Low-Level Optimization for GPUs CODE(0x5649a7648030)eaf ] ] [ Model-System Co-design [ Mixture of Experts CODE(0x5649a7648030)eaf ] [ Fine-tuning with System Optimizations CODE(0x5649a7648030)eaf ] [ Mixed Precision Training CODE(0x5649a7648030)eaf ] [ Efficient Model Inference on Optimized Systems CODE(0x5649a7648030)eaf ] [ Model Compression with FlexGen and System Optimizations CODE(0x5649a7648030)eaf ] ] ]
Figure 1: Efficient Foundation Models Overview
1 Introduction
The training and inference processes of foundation models are computationally intensive and require vast amounts of data (Wan et al., 2023; Wang et al., 2024a; Tao et al., 2024; Xiong et al., 2024a). Foundation models include multimodal models capable of handling diverse data types such as text, images, and audio. According to statistics (Bubeck et al., 2023), training models like GPT-4 can take 4–7 months, requiring thousands of GPUs, and results in models with billions of parameters. Such a high demand for computational resources, combined with the large size of the resulting models, makes a barrier for many researchers and organizations who want to train large language models but lack the financial resources to cover the computation expense. It also poses challenges for individuals who want to deploy such models on personal computers or mobile devices but lack sufficient local computational resources. Given these challenges, we anticipate that efficient design for foundation models will become a major focus in the next wave of AI advancements.
Efficient design for foundation models can generally be approached from three key perspectives: Model Design, System Design, and Model-System Co-Design.
2 Model Design for Efficient Foundation Models
2.1 Quantization
Quantization is a technique that reduces the numerical precision of model parameters and activations to improve computational efficiency. By representing values with lower-bit data types, it significantly reduces memory usage and accelerates inference. This is particularly important for deploying large-scale foundation models on resource-constrained hardware. In general, quantization methods can be categorized into bit-width based quantization and method-based quantization. Bit-width based quantization lowers precision using formats like INT8, FP8, and even lower-bit representations such as 4-bit and 2-bit. These methods apply scaling factors, Hessian-aware adjustments, and structured weight transformations to minimize accuracy loss while improving efficiency. Floating-point quantization, such as FP8, further enhances adaptability by handling activations with large dynamic ranges. Method-based quantization includes Post-Training Quantization and Quantization-Aware Training. PTQ quantizes a pre-trained model without modifying its weights, enabling fast deployment with minimal computational cost. In contrast, QAT incorporates quantization into the training process, allowing models to adapt to lower precision. Additionally, mixed-precision quantization assigns different bit-widths to different layers - it dynamically assigns bit-widths per layer to balance efficiency and accuracy, leveraging advanced techniques like adaptive scaling, reordering transformations, and structured low-rank decomposition.
By leveraging these techniques, quantization enables efficient large-model training and inference while preserving performance. The following sections explore these strategies in more detail.
2.1.1 Bit-width Based Quantization
Bit-width based quantization reduces numerical precision in model parameters and activations, optimizing memory usage and computational efficiency. 8-bit quantization methods, such as GPTQ (Frantar and Alistarh, 2022) and SmoothQuant (Xiao et al., 2023), primarily rely on fixed-point representations like INT8 and FP8 to maintain numerical stability. The standard quantization function follows $q(x)=\text{round}(x/s)$, where $s$ is a learned scale factor. FP8 methods such as F8Net (Dettmers et al., 2023) introduce fixed-point multiplication, eliminating INT32 overhead via dyadic scaling $q(x)=\text{round}(x\cdot 2^{-p})$, where $p$ dynamically minimizes quantization error. Compared to INT8, FP8 formats such as E5M2 and E4M3 have a wider dynamic range, adapting better to activations with large outliers, leading to improved numerical stability in LLMs and Vision Transformers (Shen et al., 2024).
Low-bit quantization (4-bit) approaches such as GPTQ (Frantar et al., 2022b) and AWQ (Huang et al., 2024) apply per-layer Hessian-aware weight quantization to minimize accuracy degradation, solving $\min_{\hat{W}}\|W-\hat{W}\|_{H}^{2}$, where $H$ is the Hessian capturing weight sensitivity. Atom (Zhao et al., 2024b) further refines this with mixed-precision weight-activation quantization, applying per-cluster scale factors $Q(w)=\text{round}(w/s_{h})$ for high-magnitude weights and $Q(w)=\text{round}(w/s_{l})$ for low-magnitude ones. Additionally, DilateQuant (Zhang et al., 2023d) introduces weight dilation, defining a dilation factor $s_{\text{dilate}}$ such that $\tilde{W}=(W/\max|W|)\cdot s_{\text{dilate}}$, ensuring stable quantization across weight ranges. Hybrid 4-bit formats, such as NF4 and FP4, employ logarithmic distributions to preserve high-magnitude weight precision while compressing low-importance activations.
For extreme quantization (2-bit and 1-bit), BiDM (Zheng et al., 2024b) employs Timestep-friendly Binary Structure (TBS), applying the binary activation function $A_{b}=\text{sign}(A)$ with a straight-through estimator (STE) for gradient propagation, $\frac{\partial L}{\partial A}\approx\frac{\partial L}{\partial A_{b}}$. RotateKV (Su et al., 2025a) mitigates degradation by applying rotation-based quantization transformations $\tilde{W}=RWR^{T}$, where $R$ is an adaptive rotation matrix computed via Fast Walsh-Hadamard Transform (FWHT). A generalized formula for extreme quantization using mixed-binarization strategies follows:
$q(x)=\arg\min_{q\in Q}\|x-q\|^{2}+\lambda\sum_{j}C(q_{j}),$ (1)
where $C(q_{j})$ accounts for quantization constraints across different bit-widths. This formulation captures a wide range of low-bit strategies.
Table 1: Quantization Methods for Foundation Models
Model Type Quantization Method 1-bit 2-bit 4-bit 6-bit 8-bit BFLOAT INT/FP
LLMs GPTQ INT
LLMs AWQ INT
LLMs QRazor INT
LLMs SmoothQuant FP
LLMs ZeroQuant Mixed
LLMs OmniQuant FP
LLMs Quarot FP
LLMs SpinQuant Mixed
Diffusion BiDM INT
Diffusion Q-Diffusion FP
Diffusion TFMQ-DM INT
Diffusion MPQ-DM Mixed
Diffusion PTQD Mixed
Diffusion Q-DM INT
Diffusion DilateQuant FP
2.1.2 Method-Based Quantization
Post-Training Quantization (PTQ) methods quantize models after training, offering fast deployment without retraining. GPTQ (Frantar et al., 2023) leverages second-order Hessian-aware loss minimization, solving $\min_{\hat{W}}\|W-\hat{W}\|_{H}^{2}$. PTQD (Wu et al., 2023) introduces variance calibration, modeling quantization noise as $\Delta_{\text{quant}}=\hat{\mu}-\mu+\sigma\epsilon$, where $\mu$ and $\sigma$ are activation statistics, and $\epsilon$ is Gaussian noise. Reorder-based PTQ (RPTQ) (Yuan et al., 2023) refines layer-wise quantization by applying reordering matrices $P$ and $Q$, reformulating weights as $\tilde{W}=PWQ^{-1}$, effectively reducing outlier sensitivity. FP8-based PTQ (Shen et al., 2024) exploits dynamic scaling, adjusting $e_{\text{scale}}$ in $Q(x)=\text{round}(x\cdot 2^{e_{\text{scale}}})$ per layer, ensuring uniform activation distribution across Transformer layers.
Quantization-Aware Training (QAT) incorporates quantization during training, allowing adaptation to low-bit constraints. EfficientDM (Yang et al., 2023) fine-tunes diffusion models with quantization-aware low-rank adapters (QALoRA), optimizing $Y=\text{QU}(X,s_{x})\cdot\text{QU}(W+BA,s_{w})$, where $B$ and $A$ are low-rank matrices preserving accuracy. DoReFa-Net (Zhou et al., 2016) utilizes logarithmic quantization $q(x)=2^{\lfloor\log_{2}(|x|)\rfloor}\text{sign}(x)$, preventing information loss at ultra-low bit widths. Q-Diffusion (Li et al., 2023b) employs per-timestep mixed-precision strategies, dynamically adjusting $s_{t}$ in $Q(X_{t})=\text{round}(X_{t}/s_{t})\cdot s_{t}$, stabilizing diffusion process quantization.
Mixed-Precision Quantization dynamically adjusts bit-width allocations per-layer to balance efficiency and accuracy. MPQ-DM (Wu et al., 2023) introduces per-layer mixed-precision assignment via $q(x,b)=\text{round}(x/s_{b})$, where $s_{b}$ is optimized for different bit-widths. Quant-LLM (Yao et al., 2023) employs FP6-centric quantization, formulating an optimal bit allocation problem:
$\min_{s_{b}}\sum_{i}\left(x_{i}-Q(x_{i},b)\right)^{2}+\lambda\sum_{j}C(b_{j}),$ (2)
where $C(b_{j})$ represents computational cost constraints. DilateQuant (Zhang et al., 2023d) extends weight dilation techniques, ensuring adaptive scaling across Transformer layers. Adaptive Quantization strategies such as RotateKV (Su et al., 2025a) compress KV caches via rotation-based transformations $\tilde{K}=RK,\quad\tilde{V}=RV$, minimizing information loss. PackQViT (Dong et al., 2023) optimizes vision models using logarithmic quantization $Q(x)=2^{\lfloor\log_{2}(x)\rfloor}$, further improving bit precision trade-offs.
Series GPU Model FP64 FP32 FP16 BF16 INT8 FP8 TF32
H Series H100
H800
GH200
A Series A100
A800
A30
A10
L Series L40
L40s
L4
RTX RTX 6000 Ada
RTX A6000
RTX A5000
GeForce RTX 4090 Limited
RTX 4080 Limited
RTX 3090
Jetson Jetson Orin NX
Jetson Xavier NX
Table 2: NVIDIA GPU Precision Support Summary
2.2 Knowledge Distillation
Knowledge Distillation (KD) has emerged as a fundamental strategy in optimizing large language models (LLMs) for deployment in environments with limited computational resources. As LLMs increase in complexity and size to achieve high performance across diverse language tasks, distillation techniques provide a means to compress these models while retaining essential capabilities. In the context of LLMs, KD is divided into Soft Knowledge Distillation and Hard Knowledge Distillation, each playing a unique role in balancing model size, accuracy, and efficiency. Soft distillation enhances LLM generalization by transferring nuanced probability distributions, while hard distillation prioritizes specific label accuracy and simplifies training. These techniques are pivotal in the ongoing development of compact yet capable LLMs, allowing them to retain high performance even when scaled down significantly.
2.2.1 Soft Knowledge Distillation
Soft distillation leverages the full probability distribution generated by an LLM teacher model, capturing subtleties in the relationship between various output classes. For LLMs, this means the student model not only learns the correct answers but also internalizes the nuances in the teacher’s responses to a broad array of inputs. By transferring these probability distributions, soft distillation helps the student model mimic the teacher’s language understanding, sentiment detection, and reasoning across tasks. Applying temperature scaling to soften the teacher’s outputs, the student LLM can absorb richer linguistic and semantic representations, making it particularly effective for LLMs intended for general-purpose language understanding. As a result, soft distillation is integral for developing efficient LLMs that retain robust performance across diverse tasks.
Probability Distribution Probability distribution methods involve transferring the full output probabilities from the teacher model to the student. This helps the student capture nuances in class relationships, learning not only the correct answers but also the likelihood of various alternatives. The choice of divergence function of probability distribution in KD shapes the way the student model aligns with the teacher’s output distribution, impacting both diversity and quality of generated content. A commonly used approach is forward Kullback-Leibler (KL) divergence, which minimizes the divergence by aligning the student distribution to cover all modes of the teacher distribution, as seen in work by (Hinton et al., 2015), (Romero et al., 2015) and (Zagoruyko and Komodakis, 2017). This method ensures that the student model captures the full spectrum of the teacher’s probability distribution, even for tokens assigned low probability by the teacher. However, this "mode-covering" approach may lead the student to overestimate these low-probability regions, potentially resulting in hallucinations or lower-quality generations. Alternatively, reverse KL divergence focuses on aligning the student with only the high-probability regions of the teacher distribution, as demonstrated by Gu et al. (2024b). However, reverse KL can come at the expense of diversity, often producing fewer output variations. In Baby LLaMA, temperature scaling is applied to the logits of teacher models (e.g., GPT-2 and LLaMA) to provide smoother output distributions for distillation (Timiryasov et al., 2023). This allows the student model to retain the nuances of the teacher’s probabilistic predictions, especially in low-data regimes. The soft probabilities are controlled by the temperature parameter $T$, ensuring that the model retains fine-grained knowledge:
$p_{i}=\frac{\exp(z_{i}/T)}{\sum_{j}\exp(z_{j}/T)}.$
Expanding on this, KPTD adopts a similar approach, but it leverages entity definitions and domain-specific knowledge to generate the transfer set. The student model learns to match the teacher’s output distribution based on this enriched set of entities, which is crucial for maintaining up-to-date knowledge in pre-trained models (Mirzadeh et al., 2020; Gong et al., 2022; Chen et al., 2020). Following this, the authors of GKD refine this further by using token-level forward and reverse divergence, focusing on critical tokens during distillation (Jafari et al., 2021; Kim and Rush, 2016). This design ensures that the student captures vital aspects of sequence-level tasks, such as machine translation and summarization, where token importance fluctuates throughout the sequence.
In parallel with previous work, in MetaICL (Min et al., 2022) and Multitask-ICT (Wang et al., 2023), soft distillation is adapted to transfer in-context learning (ICL) abilities from large models to smaller ones, helping the student generalize across tasks. These models use few-shot learning to transfer the multitask capabilities of large LLMs, effectively allowing the student to leverage the soft output distributions for task generalization. After this, the work from Zhao et al. (2024a) further enhances this process by introducing a multi-stage collaborative approach, where the student generates refined pseudolabels to improve the next stage of distillation, pushing the boundaries of semi-supervised learning through iteration.
Above all, the soft distillation loss can be expressed as:
$\mathcal{L}_{\text{distill}}=\lambda_{\text{fwd}}\sum_{i=1}^{N}\alpha_{i}\cdot D_{\text{KL}}\left(p_{\text{teacher},i}\|p_{\text{student},i}\right)+\lambda_{\text{rev}}\sum_{i=1}^{N}\beta_{i}\cdot D_{\text{KL}}\left(p_{\text{student},i}\|p_{\text{teacher},i}\right),$
$D_{\text{KL}}(p\|q)=\sum_{j}p_{j}\log\frac{p_{j}}{q_{j}}$
where:
  • $p_{\text{teacher},i}$ and $p_{\text{student},i}$ are the teacher and student probability distributions for token $i$, respectively.
  • $\lambda_{\text{fwd}}$ and $\lambda_{\text{rev}}$ are tunable weights balancing forward and reverse KL components.
  • $\alpha_{i}$ and $\beta_{i}$ are token-specific importance weights, which could depend on token-level saliency, attention scores, or domain-specific criteria.
  • $N$ is the total number of tokens in the sequence.
2.2.2 Hard Knowledge Distillation
Hard distillation simplifies the training process by focusing solely on the teacher model’s “hard” labels, or binary outputs, for each task. In the case of LLMs, hard distillation emphasizes specific target outcomes over nuanced inter-class information, which can be advantageous for task-focused LLMs where interpretability and task accuracy are prioritized. This method is often used in classification tasks where the LLM’s primary purpose is to deliver accurate, targeted answers without requiring complex linguistic or semantic generalization. Although this approach may limit the student’s understanding of intricate language structures, it accelerates training and achieves effective LLM compression, making it a suitable choice for specialized language tasks where efficiency and specific label accuracy are crucial.
Program-aided Distillation (PaD) addresses the challenge of faulty reasoning in distillation by introducing programmatic reasoning (Shridhar et al., 2023). Synthetic programs are generated to ensure correct reasoning steps, and these programs are automatically checked by an interpreter before being used in distillation. This reduces faulty reasoning in student models, ensuring that the correct reasoning paths are learned alongside the correct outputs. DynaBERT applies hard knowledge distillation in dynamic model compression, progressively decreasing model size while ensuring that smaller models retain the performance of the teacher (Hou et al., 2020). By combining hard distillation with a reduction in both model width and depth, DynaBERT balances efficiency with accuracy. Similarly, LaMini-LM focuses on transferring instruction-following abilities to smaller models by using hard distillation on a curated instruction set, optimizing the student for task-specific performance (Wu et al., 2024). Zephyr integrates hard knowledge distillation with Direct Preference Optimization (dDPO) (Tunstall et al., 2023). In this approach, the student learns from both hard decisions and user preferences, blending rigid output labels with nuanced optimization for user alignment. This hybrid approach is particularly effective in conversational AI, where both categorical responses and preference-based feedback are important for generating relevant and engaging responses.
Above all, the hard distillation process can be formalized as :
$\mathcal{L}_{\text{hard}}=-\sum_{i=1}^{N}\sum_{j=1}^{C}y_{i,j}^{\text{teacher}}\cdot\log p_{i,j}^{\text{student}},$
where:
  • $y_{i,j}^{\text{teacher}}\in\{0,1\}$ is the one-hot encoded hard label for token $i$ and class $j$, generated by the teacher model.
  • $p_{i,j}^{\text{student}}$ is the predicted probability of the student model for token $i$ and class $j$, computed using a softmax over the student logits:
    $p_{i,j}^{\text{student}}=\frac{\exp(z_{i,j}^{\text{student}})}{\sum_{k=1}^{C}\exp(z_{i,k}^{\text{student}})},$
    where $z_{i,j}^{\text{student}}$ is the logit for class $j$ at token $i$.
  • $N$ is the total number of tokens in the dataset.
  • $C$ is the number of classes in the classification task.
2.3 Pruning
In addition to quantization and knowledge distillation, pruning is another efficient technique for reducing computational costs in LLMs. While quantization focuses on hardware optimization and knowledge distillation on knowledge transfer in neural networks, pruning concentrates on the model’s structural efficiency. Specifically, pruning in machine learning involves removing less important connections in the neural network while retaining the important ones, making the model more memory-efficient and faster.
Pruning methods can be broadly categorized based on their granularity into unstructured pruning and structured pruning. Structured pruning removes entire neurons or larger components of the neural network. In contrast, unstructured pruning refers to the removal of individual connections, as seen in the conventional pruning methods studied by Han et al. (2015). Additionally, semi-structured pruning like N:M sparsity (Mishra et al., 2021), is generally considered a specialized form of unstructured pruning. In general, structured pruning usually yields better inference speed improvements compared to unstructured pruning, due to its hardware-friendly design.
Pruning methods for LLMs typically remove weights based either on (1) the loss function or (2) the layer output.
(1) Loss-based. The weights are usually removed by computing metric such as gradients with respect to the training or validation loss. In some cases, pruning is performed directly through optimization techniques that minimize the model’s loss after weight removal.
(2) Layer output-based. Alternatively, other pruning methods for LLMs aim to find the following masks in a layer-wise pattern. The pruning quality is usually measured by the $\ell_{2}$-norm between the output, for given specific inputs $X_{\ell}$, of the uncompressed layer and that of the compressed one. More formally, consider a layer $\ell$ with weight matrix $W_{\ell}\in\mathbb{R}^{p}$. The objective is to find the binary mask $M_{\ell}\in[0,1]^{p}$ and possibly an updated weight matrix $\hat{W}_{\ell}$ so that the following objective is minimized,
$\operatorname*{arg\,min}_{M_{\ell},\hat{W}_{\ell}}\|W_{\ell}X_{\ell}-(\hat{W}_{\ell}\odot M_{\ell})X_{\ell}\|^{2}_{2}.$ (3)
Here, $\odot$ denotes element-wise multiplication.
2.3.1 Unstructured Pruning
In this section we consider unstructured pruning and semi-structured pruning methods like N:M sparsity.
Saliency-Based Pruning.
There are several works motivated by classical post-training methods, including Optimal Brain Damage (OBD; LeCun et al. (1989)) and Optional Brain Surgeon (OBS; Hassibi and Stork (1992)). In the original OBS framework, the removal of connection at index $m$ [^1] Note that we use subscript $S_{ij}$ and $S_{m}$ interchangeably. $S_{ij}$ emphasizes the weight within a specific matrix, while $S_{m}$ emphasizes the weight across the entire neural network. with weight $w_{m}$ is selected based on the saliency metric
$S_{m}=\frac{W_{m}^{2}}{[H^{-1}]_{mm}},$
where $H$ is the Hessian matrix.
After the weight removal, all the remaining weights will be updated with the following formula:
$\delta_{m}=-\frac{w_{m}}{[H^{-1}]_{mm}}[H^{-1}]_{:,m}$
to compensate for the removal of the weight at index $m$.
Note that under the formulation of Equation 3 for an affine transformation, each row can be computed independently. Specifically, we can iteratively apply OBS to different rows by (1) identifying the location $m$ that results in the smallest loss increase, (2) constructing the updated mask $M-\{m\}$, and (3) updating all remaining weights responding to $M-\{m\}$ upon the removal of mask $\{m\}$.
Consider a weight matrix $W_{\ell}\in\mathbb{R}^{d_{row}\times d_{col}}$, SparseGPT (Frantar and Alistarh, 2023) extends the idea of OBS by updating only a subset of remaining weights $H\subseteq M-\{m\}$ to enable faster computation of the Hessian matrix. More specifically, SparseGPT processes columns sequentially from left to right using a sequence of Hessian matrices, which can be efficiently computed using Gaussian elimination. SparseGPT only uses the weights to the right of the pruned weight as subset to compensate the loss induced by its removal. The computational cost of SparseGPT is hence $\Theta(d_{hidden}^{3})$ compared to $\Theta(d_{hidden}^{4})$ required by exact construction in Transformers with hidden dimension $d_{hidden}$.
SparseGPT is also compatible with N:M sparsity by choosing N smallest weights that incurs the lowest errors in every block of M. Additionally, it also supports quantization method like GPTQ (Frantar et al., 2022a).
Wanda (Sun et al., 2023a) proposes a simple and effective pruning metric:
$S_{ij}=|W_{ij}|\cdot\|X_{j}\|_{2},$
where $W$ is a weight matrix, and $X_{j}$ is the $j$-th feature normalized by batch and sequence length dimensions. Wanda shows the connection between its pruning criterion and those of OBS and SparseGPT through a diagonal approximation of SparseGPT’s saliency score:
$S_{ij}=\left[\frac{|W|^{2}}{\text{diag}((XX^{T}+\lambda I)^{-1})}\right]_{ij}.$
Applying the diagonal approximation with $\lambda=0$, the metric can be simplified to:
$\left[\frac{|W|^{2}}{\text{diag}((XX^{T}+\lambda I)^{-1})}\right]_{ij}\stackrel{{\scriptstyle\lambda=0}}{{\approx}}\left[\frac{|W|^{2}}{(\text{diag}(XX^{T}))^{-1}}\right]_{ij}=(|W_{ij}|\cdot\|X_{j}\|_{2})^{2}.$ (4)
Compared to the complexity $\Theta(d_{hidden}^{3})$ of SparseGPT, Wanda enjoys the complexity of $\Theta(d_{hidden}^{2})$. Additionally, Wanda does not need to update the weights.
Empirical evaluations on zero-shot tasks and perplexity on LlaMA (Touvron et al., 2023a) models show the competitive performance of Wanda compared to SparseGPT. Wanda is also compatible with semi-structured N:M pruning and LoRA fine-tuning (Hu et al., 2021b).
Improved Saliency Criterion (ISC; (Shao et al., 2024)) proposed to combine criterion from both OBD and OBS to define a better saliency score:
$S_{m}=W_{m}^{2}\left(H_{mm}+\frac{1}{[H^{-1}]_{mm}}\right).$
Moreover, ISC also explores the effects of the sparsity ratio allocations between layers through sensitivity. The sensitivity is calculated by the average trace of the Hessian matrix. Experiments conducted on the LlaMA model family and Baichuan models show its strong performance and compatibility with quantization methods.
D-Pruner (Zhang et al., 2024c) studies domain-specific compression for LLMs. It uses the formulation of OBS for general knowledge and while introducing additional regularization term tailored to domain-specific knowledge. With the help of both general and domain-specific calibration datasets, D-Pruner produces a pruned model that strikes a balance between both generality and specificity.
RIA (Zhang et al., 2024d) incorporates relative importance into Wanda’s saliency score, making the final saliency score:
$S_{ij}=\left(\frac{|W_{ij}|}{\sum_{i}|W_{ij}|}+\frac{|W_{ij}|}{\sum_{j}|W_{ij}|}\right)\cdot(\|X_{i}\|_{2})^{\alpha}.$ (5)
Moreover, RIA introduces an optimization specifically for N:M sparsity through channel permutation. This enables the retention of important weights within the same block, improving the performance of pruned models.
Pruner-Zero (Dong et al., 2024) formulates pruning as Symbolic Regression (SR) problem and use genetic programming (GP) to solve it. It directly searches pruning saliency score using LLM statistics as input features, including activations ($X$), gradients ($G$), and weights ($W$). The search space includes primitive operations include unary and binary ones. Pruner-zero iteratively generates and refines pruning metrics via genetic programming procedures including tournament selection, subtree crossover, and node mutation.
There are other studies focus more on deployment and evaluation rather than solely on algorithm design. For example, Prune-And-Tune (Syed et al., 2023) demonstrates that fine-tuning can further improve the performance of SparseGPT for OPT model family, particularly at high sparsity levels. Shamrai (2024) specifically evaluates SparseGPT and Wanda specifically for the Ukrainian language.
Optimization Based Pruning.
Another line of research aims at finding masks through optimization.
BESA (Xu et al., 2024) efficiently learns block-wise sparsity ratios for each model component/block, such as Transformer blocks and feed-forward network (FFN) blocks, through minimizing the reconstruction error. The training process uses a standard Straight-through estimator (STE) to handle the non-differentiability of the masks. The sparsity is enforced through a $\ell_{2}$ regularization term during the training process.
Once the sparsity ratios are learned, Wanda saliency score will be used to prune unimportant weights within each block. Additionally, BESA is also compatible with quantization techniques.
2.3.2 Structured Pruning
Saliency-based Pruning.
Ma et al. (2023) highlights that task-specific model compression through knowledge distillation can be time-consuming in some cases, which requires up to 3.5 days on 4 GPUs for a 20GB corpus. In contrast, pruning can significantly compress model size on a structural level, which requires up to 3 hours on 1 GPU for a 50MB corpus after pruning (Wang et al., 2024b).
One famous method of pruning in LLMs is LLMPrunner (Ma et al., 2023). Initially, a dependency tree or graph is established based on structural dependency in LLMs, linking neurons based on dependency relationships. After building the dependency graph, the pruner then analyzes the importance of each link in the graph and trims those links with less importance to save computational resources and accelerate inference. LLMPrunner uses multiple methods (element-wise, vector-wise, group) to evaluate the importance of coupled structures in the dependency graph, removing less important structures. It also offers reliable recovery methods such as Low-rank Approximation to recover links discarded by mistake. Empirical results show that LLMPrunner can achieve 20% parameter reduction while retaining 90% of the original performance, showcasing the effectiveness of pruning in maintaining model efficiency while reducing computational resources by removing redundant elements in the network.
FLAP (An et al., 2024) measures the importance of different channels using fluctuation-based metrics with calibration data. It then normalizes these metrics to to ensure a consistent comparison of scores across different layers and modules. Additionally, FLAP introduces a technique called baseline bias compensation to deal with the damage caused by removing components.
Zhang et al. (2024a) proposes a structured pruning method that groups kernels and features independently and evaluates their importance by second-order Taylor expansion. To reduce computation, they approximate the second order term with squared gradient. Specifically, for kernel group $C$, the saliency score $S_{C}$ is computed as:
$\left|\sum_{c\in C}\frac{\partial\mathcal{L}}{\partial c}c-\frac{1}{2}\left(\frac{\partial^{2}\mathcal{L}}{\partial c^{2}}c^{2}\right)\right|\approx\left|\sum_{c\in C}\frac{\partial\mathcal{L}}{\partial c}c-\frac{1}{2}\left(\frac{\partial\mathcal{L}}{\partial c}c\right)^{2}\right|=S_{C},$
where $\mathcal{L}$ denotes the loss function and $c$ represents the parameters in the group $C$.
Similarly, for features $f$, the corresponding saliency score of an affine transformation is
$S_{f}=\sum_{C}\left|\sum_{c\in C[:,f]\cup C[f,:]}\frac{\partial\mathcal{L}}{\partial c}c-\frac{1}{2}\left(\frac{\partial\mathcal{L}}{\partial c}c\right)^{2}\right|.$
Optimization-based Pruning.
ShearedLlaMA (Xia et al., 2023b) formulates the pruning problem as a constrained optimization problem using Lagrange multiplier. It learns pruning masks at varying granularities, from coarse ones like entire layers and hidden dimensions to fine-grained ones like attention heads and intermediate dimensions.
Encoder-Decoder Architecture.
NASH (Ko et al., 2023) specifically studies structured pruning for encoder-decoder Transformer architecture. Inspired by empirical findings, NASH proposes to shorten, i.e., reduce depth, for decoders and narrow, i.e., reduce width, for encoders. To achieve better performance, NASH additionally applies hidden states distillation by adding a Kullback–Leibler (KL) divergence loss term.
Coupled with PEFT.
Light-PEFT (Gu et al., 2024a) and V-PETL (Xin et al., ) aims to reduce the costs of fine-tuning at multiple granularities. At a coarse level, it measures the importance of PEFT modules through the change of PEFT modules on the original module output. At a finer level, it measures the rank of each PEFT modules by summing the first-order Taylor expansion saliency score of all parameters of that rank:
$S_{i,j}=\left|\frac{\partial\mathcal{L}}{\partial W_{i,j}}W_{i,j}\right|.$
Similarly, LoRAPrune (Zhang et al., 2024b) uses the same Taylor expansion saliency score for measuring group importance. It inserts learnable LoRA matrices during downstream task adaption, circumventing the needs of computing the gradients of the entire weight matrix.
2.3.3 Others
Infra.
FlashLLM (Xia et al., 2023a) offers specific infrastructure support for unstructured pruning by utilizing Tensor Cores rather than SIMT Cores to enable faster inference.
3 System Design for Efficient Foundation Models
Large Foundation Models face significant challenges in managing computational and memory costs during autoregressive generation. Effective system design leverages optimizations like Key-Value (KV) caching, token reduction, and precision scaling to balance efficiency and accuracy. These methods are grounded in system design that formalize the trade-offs between model performance and resource utilization.
3.1 KV Cache Design
Input PromptGenerate Query (Q)Check KV Cache Update KV Cache $\mathbf{K}_{\text{low}}=\mathbf{U}\mathbf{K},\,\mathbf{V}_{\text{low}}=\mathbf{V}\mathbf{V}^{\top}$ Compute Attention $\hat{\mathbf{X}}_{t}=\text{softmax}\left(\frac{\mathbf{Q}_{t}\mathbf{K}_{1:t}^{\top}}{\sqrt{d_{k}}}+\mathbf{B}_{t}\right)\mathbf{V}_{1:t}$ Generate Output Token Optimize KV Cache $\min_{\mathbf{K}_{\text{low}},\mathbf{V}_{\text{low}}}\|\mathbf{K}-\mathbf{K}_{\text{low}}\|_{F}^{2}+\|\mathbf{V}-\mathbf{V}_{\text{low}}\|_{F}^{2}+\lambda\mathcal{C}(\mathbf{K}_{\text{low}},\mathbf{V}_{\text{low}})$
Figure 2: KV Cache Design Pipeline
KV caching is a cornerstone of transformer-based LLMs, significantly reducing computational overhead by reusing previously computed key ($\mathbf{K}$) and value ($\mathbf{V}$) matrices. During autoregressive generation, the self-attention mechanism computes the output for a sequence of tokens by leveraging the precomputed cache. Let $\mathbf{Q}_{t}$, $\mathbf{K}_{1:t}$, and $\mathbf{V}_{1:t}$ denote the query, key, and value matrices at time $t$. The attention mechanism scales the query-key interaction by the inverse square root of the key dimension $d_{k}$, applies positional biases $\mathbf{B}_{t}$, and generates the final representation $\hat{\mathbf{X}}_{t}=\text{softmax}((\mathbf{Q}_{t}\mathbf{K}_{1:t}^{\top})/\sqrt{d_{k}}+\mathbf{B}_{t})\mathbf{V}_{1:t}$. To improve memory efficiency, advanced KV cache compression methods use low-rank approximations, where keys and values are projected into lower-dimensional spaces using learned projection matrices $\mathbf{U}$ and $\mathbf{V}$. Specifically, the compressed representations are given by $\mathbf{K}_{\text{low}}=\mathbf{U}\mathbf{K}$ and $\mathbf{V}_{\text{low}}=\mathbf{V}\mathbf{V}^{\top}$, where $r\ll d_{k}$ controls the rank of the approximation. Query-aware sparsity further optimizes KV caching by dynamically identifying and updating only the most relevant tokens using a sparsity mask $\mathbf{M}_{t}$, defined as the top-$k$ entries of the scaled query-key scores. This hierarchical approach combines memory reduction and computational efficiency while maintaining the fidelity of attention outputs.
The general optimization for KV cache design balances memory usage, computational cost, and attention fidelity, expressed as:
$\min_{\mathbf{K}_{\text{low}},\mathbf{V}_{\text{low}}}\|\mathbf{K}-\mathbf{K}_{\text{low}}\|_{F}^{2}+\|\mathbf{V}-\mathbf{V}_{\text{low}}\|_{F}^{2}+\lambda\cdot\mathcal{C}(\mathbf{K}_{\text{low}},\mathbf{V}_{\text{low}}),$
where $\mathcal{C}(\mathbf{K}_{\text{low}},\mathbf{V}_{\text{low}})$ models the memory and computational costs of the compressed representations (Xiao et al., 2024b; Xiong et al., 2024b).
Autoregressive decoding in LLMs requires KV cache, which stores intermediate outputs for all tokens, creating significant memory overhead. Some adaptive KV cache compression techniques will adjust the precision of stored tokens based on their importance. Let $K,V\in\mathbb{R}^{T\times d}$ represent the key and value matrices. The optimization for adaptive precision is:
$\min_{\hat{K},\hat{V}}\|K-\hat{K}\|_{F}^{2}+\|V-\hat{V}\|_{F}^{2}+\gamma\cdot\sum_{t=1}^{T}P_{t}\cdot\mathcal{C}(P_{t}),$
where $\hat{K}$ and $\hat{V}$ are the compressed representations, $P_{t}$ denotes the precision level for token $t$, and $\mathcal{C}(P_{t})$ quantifies the cost of using a specific precision. Precision levels can be dynamically assigned as $P_{t}=\text{argmax}_{p\in\{\text{FP32},\text{FP16},\text{FP8},\text{INT8},etc\}}\Big{\{}S_{t}\cdot\text{Utility}(p)\Big{\}}$, where $S_{t}$ is the token saliency score and $\text{Utility}(p)$ reflects the precision’s effectiveness. VL-Cache achieves up to 90% memory reduction with a 7$\times$ improvement in decoding latency (Tu et al., 2024a).
3.2 Token Reduction
Token reduction focuses on minimizing the number of tokens processed during training and inference while retaining essential information for accurate predictions. In transformer models, tokens often exhibit redundancy, particularly in the deeper layers where contextualized embeddings converge. Let $T\in\mathbb{R}^{n\times d}$ denote the token embeddings, where $n$ is the token count and $d$ is the embedding dimension. Token pruning techniques dynamically identify and remove less important tokens based on learned importance scores or predefined heuristics. Importance scores can be computed using attention weights, token magnitudes, or cumulative relevance across layers. For example, attention-based pruning ranks tokens by their cumulative importance $S_{i}=\sum_{l=1}^{L}\sum_{j=1}^{n}A_{ij}^{(l)}$, where $A^{(l)}$ represents the attention matrix at layer $l$. Tokens with the lowest scores are removed, and their information is redistributed to the remaining tokens. Alternatively, token merging methods combine similar tokens, reducing the total count while preserving overall representation. In methods like Token Merging (ToMe), merged tokens are computed as weighted combinations, where similarity scores guide the merging process. The merged embeddings $\hat{T}_{k}=\alpha T_{i}+(1-\alpha)T_{j}$, with $\alpha$ determined by cosine similarity, ensure that merged tokens retain meaningful features. These approaches effectively reduce computational costs in both text and visual transformer architectures.
The optimization objective for token reduction can be formulated as:
$\min_{M,\hat{T}}\|T\odot M-\hat{T}\|_{F}^{2}+\mu\cdot\text{Cost}(M)+\nu\cdot\text{Regularization}(M),$
where $M\in\{0,1\}^{n}$ is the token retention mask, $\text{Cost}(M)$ captures the computational overhead, and $\text{Regularization}(M)$ enforces structural constraints (Bolya et al., 2023; Rao et al., 2021).
For visual LLMs, visual tokens encode rich spatial and semantic information but often contain redundancy. PyramidDrop adopts token reduction framework by computing redundancy scores for each token based on local similarity metrics (Xing et al., 2024). Similarly, Victor leverages token summarization mechanisms, where token embeddings are aggregated into compact registers to reduce computation by over 3$\times$ with minimal accuracy loss (Wen et al., 2024).
3.3 Sparsity-Aware Optimization
The QKV transformation requires masking for autoregressive prediction. The structure $QK^{\top}\odot M$ results in a lower triangular masked attention matrix of size $N\times N$, where $N$ is the processed sequence length, which can be significantly large in large-scale language models.
However, not all tokens contribute equally to the final output. Studies show that a small subset of tokens dominates attention computations, while others carry minimal impact. Optimizing attention sparsity reduces computational overhead without sacrificing performance. Several methods address this:
Token Selection and Structured Sparsity
Some Research reduces computational costs by focusing only on high-saliency tokens. Let $A\in\mathbb{R}^{T\times T}$ represent the attention matrix at a given layer, where $A_{ij}$ denotes the attention weight between tokens $i$ and $j$. Sparsity-aware optimization selects a subset of tokens $S=\{t\mid S_{t}>\tau\}$, where $S_{t}=\sum_{j=1}^{T}A_{tj}$ represents the token’s saliency score and $\tau$ is a dynamically determined threshold. The optimization problem can be expressed as:
$\max_{\rho_{l}}\sum_{l=1}^{N_{L}}\rho_{l}\cdot\sum_{t\in S}\log\Big{(}1+\frac{S_{t}}{\tau}\Big{)}-\kappa\cdot\mathcal{C}(\rho_{l}),$
where $\rho_{l}$ denotes the retention ratio at layer $l$, $\mathcal{C}(\rho_{l})$ is the computational cost associated with retaining tokens, and $\kappa$ balances performance and cost. Methods like ZipVL and $H_{2}O$ dynamically adjust $\rho_{l}$ across layers based on the attention distribution (He et al., 2024; Zhang et al., 2023c), achieving over large memory and cache savings with small accuracy degradation.
Activation and Model Weight Sparsity
Beyond attention sparsity, recent studies have explored the natural sparsity in transformer activations. It has been observed that ReLU-based transformers exhibit a high degree of activation sparsity in the feedforward network (FFN) layers (Mirzadeh et al., 2023; Sun et al., 2023c). In models like OPT-175B, over 95% of FFN activations are zero. While modern transformers favor soft activations such as GeLU and SwiGLU, replacing them with ReLU and fine-tuning has been shown to increase activation sparsity while maintaining performance (Mirzadeh et al., 2024).
Another effective sparsity technique involves weight pruning. Static weight pruning approaches, such as Wanda (Sun et al., 2023b), analyze weight importance and remove redundant connections before inference. These methods ensure a fixed sparse structure, unlike dynamic sparsity techniques that adapt weight selection based on input data. LoRA (Hu et al., 2021c) further reduces computational costs by introducing low-rank updates instead of modifying the full weight matrices. LoRA injects trainable low-rank matrices $\Delta W$ into the model:
$W^{\prime}=W+\Delta W,\quad\text{where}\quad\Delta W=AB^{\top},\quad A,B\in\mathbb{R}^{d\times r}.$
With $r\ll d$, LoRA achieves significant efficiency improvements, making fine-tuning large models more memory-efficient.
3.4 Multi-Modal Fusion
Multi-modal fusion integrates information from different modality, such as text, images, and audio, to create a richer and more comprehensive representation. By capturing relationships between modalities, fusion methods enhance model understanding and response quality. Key techniques include unified feature representations, cross-modal attention mechanisms, and efficient compression strategies. These approaches enable a combined interaction between modalities while maintaining efficiency in computation and memory usage.
3.4.1 Multi-Modal Fusion Representations
Multi-Modal Key and Value Structures
Unified KV representations reduce redundancy while capturing cross-modal dependencies. For a multi-modal input $x=\{x_{t},x_{i},x_{a}\}$ (text, image, and audio, respectively), the KV cache can be structured as:
$K=\begin{bmatrix}W_{k}^{(t)}h^{(t)}&\alpha_{ti}W_{k}^{(i)}h^{(i)}&\alpha_{ta}W_{k}^{(a)}h^{(a)}\\ \alpha_{it}W_{k}^{(t)}h^{(t)}&W_{k}^{(i)}h^{(i)}&\alpha_{ia}W_{k}^{(a)}h^{(a)}\\ \alpha_{at}W_{k}^{(t)}h^{(t)}&\alpha_{ai}W_{k}^{(i)}h^{(i)}&W_{k}^{(a)}h^{(a)}\end{bmatrix},$
where $W_{k}^{(m)}$ and $h^{(m)}$ represent the projection matrix and hidden state for modality $m$. The coefficients $\alpha_{mn}$ denote learned alignment weights between modalities $m$ and $n$, enabling efficient cross-modal interaction (Radford et al., 2021; Alayrac et al., 2022; Geng et al., 2022).
Joint Cross-Modal Attention-Based Compression
Given the computational overhead of storing full KV caches, cross-modal attention matrices $A$ can be used to guide compression. For text $T$ and image $I$, the attention matrix is:
$A_{TI}=\text{diag}(K_{T}^{\top}Q_{I})+\text{diag}(K_{I}^{\top}Q_{T}),$
where $Q$ and $K$ represent the query and key matrices, respectively. A low-rank approximation of $A_{TI}$ via Singular Value Decomposition (SVD) yields:
$A_{TI}\approx U_{TI}\Sigma_{TI}V_{TI}^{\top},\quad\Sigma_{TI}=\text{diag}(\sigma_{1},\sigma_{2},\dots,\sigma_{r}),$
where $U_{TI},\Sigma_{TI},V_{TI}$ are the top-$r$ components. This reduced-rank representation balances efficiency and accuracy (Jaegle et al., 2021; Yang et al., 2024c).
3.4.2 Cache Eviction on Multi-Modality Fusion
Cross-Modal Sensitivity-Based Eviction
Cache eviction in multi-modal systems must account for both intra-modal and cross-modal contributions. The sensitivity of the model output $y$ to a key $k_{t}^{m}$ from modality $m$ is defined as:
$S_{t}^{(m)}=\left\|\frac{\partial y}{\partial k_{t}^{m}}\right\|_{2}^{2}+\sum_{n\neq m}\beta_{mn}\left\|\frac{\partial y}{\partial k_{t}^{n}}\cdot\frac{\partial k_{t}^{n}}{\partial k_{t}^{m}}\right\|_{2}^{2},$
where $\beta_{mn}$ captures the dependency strength between modalities $m$ and $n$. Entries with the smallest $S_{t}^{(m)}$ are evicted, ensuring minimal performance degradation (He et al., 2023; Xu et al., 2025).
Hessian-Aware Cross-Modal Eviction
Second-order optimization methods are useful for assessing the importance of cache entries. The Hessian-aware importance of a key $k_{t}^{m}$ is computed as:
$\Delta\mathcal{L}=\frac{1}{2}\left(\frac{\partial\mathcal{L}}{\partial k_{t}^{m}}^{\top}H_{t}^{(m)}\frac{\partial\mathcal{L}}{\partial k_{t}^{m}}+\sum_{n\neq m}\frac{\partial\mathcal{L}}{\partial k_{t}^{m}}^{\top}H_{t}^{(mn)}\frac{\partial\mathcal{L}}{\partial k_{t}^{n}}\right),$
where $H_{t}^{(m)}$ is the Hessian for modality $m$, and $H_{t}^{(mn)}$ captures cross-modal dependencies. Cache entries with minimal $\Delta\mathcal{L}$ are prioritized for eviction, preserving critical cross-modal interactions (Alayrac et al., 2022; Ren and Zhu, 2024).
Entropy-Regularized Cross-Modal Eviction
Maintaining diversity in cache content can be achieved through entropy-based eviction. The joint entropy of a key distribution $p_{t}^{(m,n)}$ across modalities $m$ and $n$ is:
$H_{t}=-\sum_{m\in\mathcal{M}}\sum_{n\in\mathcal{M}}p_{t}^{(m,n)}\log p_{t}^{(m,n)},$
where $p_{t}^{(m,n)}$ is the normalized attention contribution of $k_{t}$ for modalities $m$ and $n$. Entries with low $H_{t}$ are evicted, promoting diversity and informativeness (Han et al., 2023; Yang et al., 2024a).
3.5 Efficient Sequence Modeling
Transformer-based LLMs has achieved state-of-the-art results across a variety of tasks (Vaswani et al., 2017; Achiam et al., 2023; Touvron et al., 2023b). The self-attention of Transformers will achieve an all-to-all information routing between tokens in a sequence, which will result in a long sequence dependent representations (Bahdanau et al., 2014). However, self-attention scales quadratically with sequence length, leading to high computational costs during training and inference(Deng et al., 2020; Zhou et al., 2024). Besides, The key-value caching during autoregressive generation further increases memory overhead (Tay et al., 2022).
To mitigate these efficiency challenges, several alternative architectures have been proposed, including linear attention, state-space models, long convolution, and hybrid methods, each offering distinct computational advantages over standard transformers.
Linear Attention
Linear attention restructures the self-attention computation to avoid the explicit quadratic complexity of softmax attention. Given standard attention:
$A=\text{softmax}\left(\frac{QK^{\top}}{\sqrt{d}}\right)V,$
where $Q,K,V\in\mathbb{R}^{N\times d}$ represent query, key, and value matrices, linear attention instead decomposes the softmax function using kernel-based approximations:
$O=\phi(Q)\left(\phi(K)^{\top}V\right),$
where $\phi(\cdot)$ is a feature transformation designed to preserve the key properties of softmax while enabling efficient factorization (Katharopoulos et al., 2020; Choromanski et al., 2021; Qin et al., 2023b). Variants of $\phi(x)$ include: - $\phi(x)=1+\text{elu}(x)$ (Katharopoulos et al., 2020). - Cosine approximations (Qin et al., 2023b). - Random feature-based mappings (Choromanski et al., 2021; Shen et al., 2023a).
Despite achieving $\mathcal{O}(Nd^{2})$ complexity, causal implementations require cumulative summation (cumsum) operations, limiting their efficiency in autoregressive settings (Hua et al., 2022). Lightning Attention (Qin et al., 2024a; b) resolves this through a structured block-wise decomposition that eliminates cumsum, maintaining efficient scaling across sequence lengths.
State Space Models
State-space models (SSMs) frame sequence modeling as a continuous dynamical system:
$h^{\prime}(t)=Ah(t)+Bx(t),$
$y(t)=Ch(t),$
where $A,B,C$ are learned transition matrices controlling the evolution of hidden states (Gu et al., 2022a). Unlike self-attention, SSMs process sequences with constant complexity, making them well-suited for long-context modeling (Gu et al., 2022b; c).
Mamba (Gu and Dao, 2024) improves upon standard SSMs by introducing selective state updates, dynamically adjusting recurrence parameters based on input tokens. The discrete recurrence formulation is given by:
$h_{t}=Ah_{t-1}+Bx_{t},\quad y_{t}=Ch_{t}.$
This design eliminates the need for extensive key-value caching, ensuring efficient autoregressive inference with linear complexity.
Long Convolution
Long convolutional models replace explicit attention mechanisms with large-kernel convolutions that capture long-range dependencies (Qin et al., 2023a; Fu et al., 2023). The core operation for a sequence $X$ with kernel $W$ is:
$O_{t}=\sum_{i=0}^{L}W_{i}X_{t-i}.$
To accelerate training, Fast Fourier Transforms (FFT) reduce computational complexity from $\mathcal{O}(N^{2})$ to $\mathcal{O}(N\log N)$. However, causal convolutions require caching all prior activations, leading to high memory usage in autoregressive applications.
Hybrid Models: Combining Attention and SSMs
While alternative architectures provide computational benefits, self-attention remains superior in capturing in-context relationships and performing few-shot generalization. Hybrid models, such as Mamba-2-Hybrid (Waleffe et al., 2024), integrate selective self-attention with state-space models, achieving a balance between efficiency and model expressivity. Empirical studies show that hybrid architectures outperform both transformers and pure SSMs in long-context tasks, efficiently handling sequences up to 128K tokens.
Model Type Time Complexity Space Complexity
Softmax Attention $\mathcal{O}(N^{2}d)$ $\mathcal{O}(N^{2})$
Linear Attention $\mathcal{O}(Nd^{2})$ $\mathcal{O}(Nd)$
State-Space Models (SSMs) $\mathcal{O}(Nd)$ $\mathcal{O}(Nd)$
Hybrid (SSM + Attention) $\mathcal{O}(Nd+Sd^{2})$ $\mathcal{O}(Nd)$
Multi-Query Attention (MQA) $\mathcal{O}(Nd^{2})$ $\mathcal{O}(Nd)$
Grouped-Query Attention (GQA) $\mathcal{O}(Nd^{2})$ $\mathcal{O}(Nd)$
Multi-Head Attention (MHA) $\mathcal{O}(N^{2}d)$ $\mathcal{O}(N^{2})$
Multi-Head Latent Attention (MLA) $\mathcal{O}(Ndr)$ $\mathcal{O}(Nr)$
Table 3: Comparison of sequence modeling approaches. $N$: sequence length, $d$: feature dimension, $S$: state dimension, $r$: reduced latent dimension in MLA.
Recent advances in sequence modeling reduce computational overhead. Linear attention employs kernel approximations to optimize efficiency, state-space models leverage structured recurrence for constant-time token generation, and long convolutional introduces FFT-based acceleration, etc. These efficient sequence processing techniques help reduce memory usage and speed up computation while keeping the model’s ability to understand long sequences. They make it possible to handle extremely long contexts in large-scale language models.
Table 4: Open Source LLM System Design Frameworks Features
Framework Key Features
TensorRT Optimized for NVIDIA hardware, KV caching support
FasterTransformer Pre-allocated KV memory for faster inference
CacheGen KV cache compression for faster inference
vLLM Memory paging with PagedAttention
FlexGen Compression-based optimization for large models
FlashInfer High-performance CUDA kernels, optimized KV caching
LlamaCPP Optimized inference for CPU, supports GGUF models
DeepSeek Fine-grained optimization for distributed training
DeepSpeed ZeRO optimization, pipeline parallelism, multi-GPU scaling
Bloom Optimized LLM inference and fine-tuning
Hugging Face Transformers Pretrained models, easy-to-use APIs, multi-platform support
ONNX Runtime Hardware-agnostic optimization, quantization support
Megatron-LM Large-scale distributed training, tensor and pipeline parallelism
Fairseq Sequence modeling, multilingual support, model compression
Colossal-AI Distributed training, memory optimization, tensor slicing
OpenFold DeepMind-inspired protein modeling, memory efficiency
Triton Highly efficient GPU kernels for custom inference optimization
Table 5: Open Source LLM System Design Frameworks Code
Framework Source Link
TensorRT https://github.com/NVIDIA/TensorRT-LLM
FasterTransformer https://github.com/NVIDIA/FasterTransformer
CacheGen https://github.com/cachegen-project
vLLM https://github.com/vllm-project/vllm
FlexGen https://github.com/FMInference/FlexGen
FlashInfer https://github.com/Dao-AILab/flashinfer
LlamaCPP https://github.com/ggerganov/llama.cpp
DeepSeek https://github.com/deepseek-project
DeepSpeed https://github.com/microsoft/DeepSpeed
Bloom https://github.com/bigscience-workshop/bloom
Hugging Face Transformers https://github.com/huggingface/transformers
ONNX Runtime https://github.com/microsoft/onnxruntime
Megatron-LM https://github.com/NVIDIA/Megatron-LM
Fairseq https://github.com/facebookresearch/fairseq
Colossal-AI https://github.com/hpcaitech/ColossalAI
OpenFold https://github.com/aqlaboratory/openfold
Triton https://github.com/openai/triton
4 Model System Co-Design for Efficient Foundation Models
4.1 Mixture of Experts (MoE)
The concept of the Mixture of experts, first introduced in (Jacobs et al., 1991; Jordan and Jacobs, 1994) has been thoroughly studied and developed through numerous subsequent works (Collobert et al., 2001; Rasmussen and Ghahramani, 2001; Eigen et al., 2014; Theis and Bethge, 2015; Aljundi et al., 2017). The advent of sparsely-gated MoE (Shazeer et al., 2017), especially in combination with transformer-based large language models (Lepikhin et al., 2020a), has revitalized this technology, which has been evolving for over three decades. The MoE framework operates on a straightforward but impactful principle: distinct components of the model, referred to as experts, are designed to specialize in specific tasks or data aspects. In this approach, only the relevant experts are activated for a given input, balancing computational efficiency with access to a vast pool of specialized expertise. This scalable and adaptable framework provides an effective solution for adhering to the scaling law, enabling increased model capacity without a proportional rise in computational costs.
The mixture of experts (MoE) approach has continued to experience significant growth, with notable advancements in 2024, including the introduction of Mixtral-8x7B (Jiang et al., 2024) and several other large-scale industrial language models such as Grok-1, DBRX, Arctic, DeepSeek-V2 (DeepSeek-AI and Liu, 2024a), and DeepSeek-V3 (DeepSeek-AI and Liu, 2024b) among others. In this work, we will use the taxonomy introduced in (Cai et al., 2024). Under this taxonomy, all the works on MoE are associated with one of three aspects: algorithmic design, system design, and application.
4.1.1 Algorithmic Design for MoE
The Mixture of Experts (MoE) framework utilizes various gating functions to determine which experts to activate for processing input data. These gating functions can be categorized into three types: sparse, dense, and soft gating.
Dense Gating
Dense gating activates all experts during each iteration. The output of the dense MoE layer can be formulated as:
$y=\sum_{i=1}^{n}G(x)_{i}.E_{i}(x)$
The $y$ is the output of MoE, $G$ is the router output for the $i$-th expert, and $E$ is the output of the $i$-th expert. The $G$ or gating function is as below:
$G_{\sigma}(x)=Softmax(x.W_{g})$
Sparse Gating
In contrast, this method activates a selected subset of experts for each input token, allowing for conditional computation. The gating function can be expressed as:
$G_{\sigma}(x)=Softmax(KeepTopK(H(x),K))$
$H(x)_{i}=(x.W_{g})_{i}$
$KeepTopK(v,k)_{i}=\begin{cases}v_{i}&v_{i}\text{ in top k elements of }v\\ -\infty&otherwise\par\end{cases}$
While the sparse gate $G_{\sigma}(x)$ significantly increases the model’s parameter capacity without proportionally raising computational costs, it can introduce a load balancing problem. This issue arises when the workload is unevenly distributed among the experts, causing some to be heavily utilized while others are rarely or never activated.
Auxiliary Loss Functions
To address this, each MoE layer incorporates an auxiliary loss function that promotes an even distribution of tokens across experts within each batch, as described in many studies (Lepikhin et al., 2020b; Jiang et al., 2024; Du and Huang, 2022; Lieber et al., 2024; Dai et al., 2024). To formulate this concept, consider a batch of queries $B=\{x_{1},x_{2},...,x_{T}\}$, comprising $T$ tokens, and $N$ experts indexed from $i=1$ to $N$ Following [28], [36], the auxiliary load balancing loss for the batch is defined as:
$L_{load-balancing}=N\sum_{i=1}^{N}D_{i}P_{i},$
$D_{i}=\frac{1}{T}\sum_{x\in B}1\{argmaxG_{\sigma}(x)=i\},$
$P_{i}=\frac{1}{T}\sum_{x\in B}G_{\sigma}(x)_{i}$
where $D_{i}$ represents the proportion of tokens distributed to expert $i$, while $P_{i}$ denotes the proportion of the gating probability assigned to expert $i$. To ensure an even distribution of the batch of tokens across the $N$ experts, the load-balancing loss function $L_{load-balancing}$ should be minimized.
Soft Gating
The allocation of experts to input tokens presents a discrete optimization challenge in sparse MoE, often requiring heuristic auxiliary losses to balance expert utilization and minimize unassigned tokens. These challenges are exacerbated in out-of-distribution scenarios such as small inference batches, novel inputs, or transfer learning. Soft MoE addresses these issues by maintaining full differentiability and using all experts for processing each input, avoiding the pitfalls of discrete expert selection.
(Puigcerver et al., 2024) introduced Soft MoE, which replaces sparse, discrete gating with a soft assignment strategy that merges tokens. This approach computes weighted averages of tokens, with weights dependent on both tokens and experts, and processes each aggregate with the corresponding expert. Experimental results in image classification show that Soft MoE improves gating function training stability and maintains balance.
4.1.2 System Design for MoE
The mixture of Experts (MoE) enhances large language models but introduces challenges due to its sparse and dynamic computational workload. GShard (Lepikhin et al., 2020a) addresses this with expert parallelism, implementing parallel gating and expert computation by dispatching partitioned local tokens with load-balancing constraints. This strategy, which extends data parallelism (Rajbhandari et al., 2020; Ren et al., 2021; Rajbhandari et al., 2021), assigns each MoE expert to a distinct device while duplicating non-expert layers across devices, enabling efficient scaling of MoE models.
Computation
Although MoE scales model parameters efficiently without increasing computational costs, it faces challenges in computational efficiency. A key issue is the uneven load distribution in expert parallelism, causing synchronization delays as the system waits for the most loaded expert to finish. Solutions include optimized gating mechanisms, expert capacity adjustments, and dynamic expert placement strategies introduced by SE-MoE (Shen et al., 2022), Tutel (Hwang et al., 2023), FlexMoE (Nie et al., 2023), and SmartMoE (Zhai et al., 2023), which aim to balance workloads across devices. FasterMoE (He et al., 2022) addresses severe imbalances with a dynamic shadowed expert strategy, replicating experts on multiple devices. These placement strategies influence both computation and communication efficiency.
Communication
In expert parallelism, the repeated All-to-All communication during forward and backward propagation in each MoE layer creates significant overhead, often limiting efficiency. This communication spans intra-node (e.g., PCIe, NVLink) and inter-node (e.g., Ethernet, Infiniband) channels, with performance affected by bandwidth heterogeneity, network topology, and collective communication algorithms. Load imbalances in MoE further increase synchronization delays.
To address these challenges, DeepSpeed-MoE (Rajbhandari et al., 2022), HetuMoE (Nie et al., 2022), and ScheMoE (Shi et al., 2024) implement hierarchical All-to-All strategies to prioritize intra-node communication and reduce inter-node exchanges. FasterMoE (He et al., 2022), TA-MoE (Chen et al., 2022), and SE-MoE (Shen et al., 2022) introduce topology-aware routing to minimize cross-node expert selection and inter-node communication. ExFlow (Yao et al., 2024) improves efficiency by leveraging expert affinity, keeping token processing within local GPUs. These strategies, combined with optimized placement of non-expert modules, enhance network efficiency and overall performance in distributed MoE systems (Rajbhandari et al., 2022; Singh et al., 2023; Wei et al., 2024).
Storage
The growing parameter sizes in MoE models strain memory capacity, a challenge already significant in dense models. Expert parallelism helps distribute experts across devices, but individual devices, especially edge devices (e.g., PCs, smartphones, IoTs), may still face storage limitations during inference.
Solutions like SE-MoE (Shen et al., 2022), Pre-gated MoE (Hwang et al., 2024), and EdgeMoE (Yi et al., 2023) address this by retaining only essential non-expert and active expert parameters in GPU High-Bandwidth Memory (HBM), while offloading inactive parameters to CPU memory or SSDs. To reduce overhead from data transfers, these methods incorporate expert selection forecasting and prefetching to overlap parameter access with computation.
4.1.3 Application for MoE
Natural Language Processing
The integration of MoE architectures into large language models (LLMs) has significantly advanced natural language understanding (NLU) and generation (NLG) tasks, including machine translation (Shazeer et al., 2017; Team and Costa-jussà, 2022), open-domain question answering (Du and Huang, 2022; Artetxe and Bhosale, 2022), code generation (Jiang et al., 2024; Wei et al., 2024; Dai et al., 2024), and mathematical problem solving (Jiang et al., 2024; DeepSeek-AI and Liu, 2024a; Dai et al., 2024). Details on incorporating MoE into LLMs are covered in algorithm design and system design. Additionally, MoE has improved LLM safety without compromising usability, as demonstrated by MoGU (Du et al., 2024), which employs dynamic routing to balance contributions between usable and safe LLMs.
Computer Vision
The success of sparsely-gated Mixture of Experts (MoE) in NLP has inspired its use in computer vision. (Riquelme et al., 2021) proposed Vision MoE (V-MoE), integrating sparsely activated MLPs into select ViT blocks. In image recognition, V-MoE matches state-of-the-art performance while reducing inference computational costs, showcasing MoE’s ability to capture distinct image semantics through specialized experts. Additionally, (Hwang et al., 2023) introduced Tutel, a scalable MoE system with dynamic parallelism and pipelining, demonstrated using SwinV2-MoE, built on Swin Transformer V2 (Liu et al., 2021b).
Multimodal Application
Multimodal models process and integrate multiple data types, often combining images and text (Baltrušaitis et al., 2019). Mixture of Experts (MoE) architectures are foundational for these models due to their ability to specialize expert layers for different modalities. A key example is LIMoE (Mustafa et al., 2022), a sparse MoE model for multimodal learning trained with contrastive loss and entropy-based regularization to address load balancing. Further advancements by (Shen et al., 2023b; Li et al., 2023c; Lin et al., 2024) have enhanced the scaling of vision-language models.
To tackle task conflicts in instruction tuning of Large Vision-Language Models (LVLMs), MoCLE (Gou et al., 2024) combines MoE with LoRA (Hu et al., 2021d) experts and a universal expert for task-specific parameter activation. Similarly, LLaVAMoLE (Chen et al., 2024) mitigates data conflicts in Multimodal Large Language Models (MLLMs) using LoRA experts for the MLP layer and a top-1 gating mechanism to refine instruction tuning.
4.2 Mixed Precision Training (MPT)
Mixed precision training integrates multiple numerical formats, such as FP32, FP16, FP8 and INT8, to optimize computational efficiency for large-scale deep learning models. By selectively assigning precision to various components of the model, it reduces memory usage and accelerates computation while preserving accuracy. A critical aspect of this approach is the dynamic assignment of precision, where components that significantly impact model performance are kept in high precision, while less critical ones are quantized. For example, let $\mathcal{F}(X;W)$ represent the output of a neural network layer with input $X$ and weights $W$. In a mixed precision setting, the weights can be decomposed as $W=W_{\text{high}}+Q_{\text{low}}$, where $W_{\text{high}}$ remains in FP32, and $Q_{\text{low}}$ is a quantized representation in FP16 or INT8. The optimal precision assignment $\mathcal{P}$ minimizes the reconstruction error $\|\mathcal{F}(X;W)-\mathcal{F}(X;\mathcal{P}(W))\|_{2}^{2}$, subject to $\mathcal{P}(W)\in\{\text{FP32},\text{FP16},\text{FP8},\text{INT8},etc\}$ (Tu et al., 2024b; Micikevicius et al., 2017; Zhao et al., 2021; Piao et al., 2022).
A notable application of precision assignment arises in attention mechanisms, where the key-value cache, denoted by $K\in\mathbb{R}^{d_{k}\times n}$ and $V\in\mathbb{R}^{d_{v}\times n}$, can benefit from dynamic scaling. For each key vector $k_{j}$, its precision can be adjusted based on an importance score. Specifically, the precision map $\mathcal{P}(k_{j})$ is designed to minimize the combined reconstruction error and alignment loss (Yang et al., 2024b). This optimization ensures that high-magnitude vectors remain in FP32, while low-magnitude ones are quantized for efficiency.
$\min_{\mathcal{P}}\mathbb{E}_{X}\bigg{[}\sum_{j=1}^{n}\Big{(}\|k_{j}-\mathcal{P}(k_{j})\|_{2}^{2}+\lambda\cdot\|\text{softmax}(K^{\top}X)-\text{softmax}(\mathcal{P}(K)^{\top}X)\|_{F}^{2}\Big{)}\bigg{]}.$
4.2.1 Computation for MPT
In mixed precision training, forward and backward passes are carefully designed to maximize efficiency without compromising stability. During the forward pass, the input to each layer, $x^{(l)}$, is multiplied by the weight matrix, $W^{(l)}$, to compute the pre-activation output $z^{(l)}=W^{(l)}x^{(l)}+b^{(l)}$, where $b^{(l)}$ is the bias term. To leverage hardware acceleration, $W^{(l)}$ is stored in FP16, while $z^{(l)}$ may temporarily reside in FP32 to prevent numerical overflow during activation. For example, when ReLU is applied, $y^{(l)}=\max(0,z^{(l)})$ is computed in FP32 to maintain accuracy. In the backward pass, gradients are computed using the chain rule, where $\nabla W^{(l)}=\delta^{(l)}(x^{(l)})^{\top}$, and $\delta^{(l)}=\sigma^{\prime}(z^{(l)})\cdot\delta^{(l+1)}$ represents the backpropagated gradient through the activation $\sigma$. To ensure numerical stability, all gradients are accumulated in FP32, even if intermediate values are stored in FP16.
To handle mixed precision more robustly, recent works propose incorporating second-order approximations for more precise weight updates (Micikevicius et al., 2017). For a general weight matrix $W$ undergoing gradient descent with mixed precision, the update rule can be expressed as:
$W_{\text{new}}=W-\eta\left[\nabla\mathcal{L}(W)+\frac{\partial^{2}\mathcal{L}}{\partial W^{2}}\cdot Q_{\text{low}}\right],$
where $\frac{\partial^{2}\mathcal{L}}{\partial W^{2}}$ is the Hessian, and $Q_{\text{low}}$ is the quantization error. This formulation allows for dynamically correcting the impact of quantized components during training CODE(0x5649a7648030)Error Correction for MPT Mixed Precision Training introduces errors that propagate through the layers of a neural network. Let $\mathcal{N}(X;W)$ represent the output of a neural operator with input $X$ and weights $W$, and $\hat{\mathcal{N}}(X;\hat{W})$ its mixed precision approximation. Using a Taylor expansion, the error propagation can be analyzed as:
$\mathcal{N}(X;W)-\hat{\mathcal{N}}(X;\hat{W})\approx J_{\mathcal{N}}(W)\cdot(W-\hat{W})+\frac{1}{2}(W-\hat{W})^{\top}H_{\mathcal{N}}(W)(W-\hat{W}),$
where $J_{\mathcal{N}}(W)$ and $H_{\mathcal{N}}(W)$ denote the Jacobian and Hessian of $\mathcal{N}$, respectively. The first-order term dominates when quantization errors are small, while the second-order term becomes significant for larger errors. To minimize error propagation, recent studies introduce regularization terms in the loss function (Tu et al., 2024b):
$\mathcal{L}_{\text{reg}}=\mathcal{L}+\beta\cdot\|J_{\mathcal{N}}(W)\cdot Q_{\text{low}}\|_{F}^{2}+\gamma\cdot\|H_{\mathcal{N}}(W)\cdot Q_{\text{low}}\|_{F}^{2},$
where $\beta$ and $\gamma$ are hyperparameters controlling the weight of first-order and second-order regularization.
4.2.2 Theoretical Analysis for MPT
A key theoretical challenge in mixed precision training is quantifying the error introduced by low-precision computations. The approximation error $\epsilon$ can be bounded as:
$\epsilon=\|\mathcal{N}(X;W)-\hat{\mathcal{N}}(X;\hat{W})\|_{F}\leq\|J_{\mathcal{N}}(W)\|_{F}\cdot\|W-\hat{W}\|_{F}+\frac{1}{2}\|H_{\mathcal{N}}(W)\|_{F}\cdot\|W-\hat{W}\|_{F}^{2}.$
This bound ensures that the total error is a function of both the quantization level and the sensitivity of the network to weight perturbations, represented by the Jacobian and Hessian norms. These guarantees are critical for designing robust precision allocation strategies and ensuring reliable model performance under mixed precision training (Das et al., 2018; Tu et al., 2024b).
4.3 Efficient Pretraining
Efficient pretraining combines model-level innovations and system-level optimizations to enable large language models (LLMs) to scale efficiently while minimizing computational and memory costs. Recent advancements focus on optimizing transformer architectures, pretraining objectives, and data strategies while aligning these improvements with hardware capabilities, such as quantization and distributed systems. This model-system co-design approach ensures that both algorithmic performance and resource utilization are optimized.
4.3.1 Efficient Pretraining Optimization
Optimizing the transformer architecture is central to efficient pretraining. GQKVA introduces grouped attention, where queries $Q$, keys $K$, and values $V$ are clustered to reduce redundancy in self-attention operations (Javadi et al., 2023). Let $Q,K,V\in\mathbb{R}^{n\times d_{k}}$ be the query, key, and value matrices. The grouped matrices $\tilde{Q},\tilde{K},\tilde{V}\in\mathbb{R}^{\kappa\times d_{k}}$ are defined by clustering with $\kappa$ groups, reducing the attention complexity from $O(n^{2}d_{k})$ to $O(\kappa^{2}d_{k})$. This optimization aligns with hardware acceleration, such as tensor cores, which are efficient for smaller matrix operations.
[…]
4.3.2 Efficient Pretraining on Distributed Systems
Continual pretraining enables general-purpose LLMs to adapt efficiently to domain-specific tasks. METRO introduces a denoising objective combined with model-generated auxiliary signals to improve robustness (Bajaj et al., 2022). Let $X_{m}$ denote a masked input sequence, and let $\mathcal{F}_{\theta}$ be the model’s prediction function. The pretraining objective is:
$\mathcal{L}_{\text{pretrain}}=\sum_{t=1}^{n}\|X_{t,\text{true}}-\mathcal{F}_{\theta}(X_{m,t})\|^{2}+\alpha\sum_{t=1}^{n}\|S_{t}-\mathcal{F}_{\theta}(X_{m,t})\|^{2},$
where $S_{t}$ are auxiliary signals generated by the model, and $\alpha$ balances the two components. This objective can be efficiently distributed across GPUs using data partitioning.
FinGPT-HPC utilizes bucketed data shuffling to optimize I/O performance during distributed training (Liu et al., 2024b). Each bucket $B_{i}$ is assigned to a GPU, where the time per training step is modeled as: $T_{\text{pretrain}}=\max\left(\frac{\mathcal{C}_{\text{comm}}}{N_{\text{GPUs}}},\mathcal{C}_{\text{comp}}\right),$
where $\mathcal{C}_{\text{comm}}$ and $\mathcal{C}_{\text{comp}}$ represent communication and computation costs. By overlapping communication and computation, this approach ensures scalability.
4.3.3 Efficient Pretraining via Sampling
Efficient pretraining also relies on prioritizing high-relevance data through task-specific sampling. Bucket Pre-training assigns each bucket $B_{i}$ a relevance weight $w_{i}$, where the sampling probability is: $p_{i}=\frac{w_{i}}{\sum_{j=1}^{m}w_{j}}$. This ensures that critical data receives more updates, reducing overfitting to irrelevant tasks (Liu et al., 2024a).
MixMAE combines mixed and masked token reconstruction objectives to enhance the diversity of token representations (Liu et al., 2023a). The total loss is:
$\mathcal{L}_{\text{MixMAE}}=\beta\cdot\mathcal{L}_{\text{mix}}+(1-\beta)\cdot\mathcal{L}_{\text{mask}}+\gamma\cdot\mathcal{L}_{\text{contrast}},$
where $\beta,\gamma\in[0,1]$ control the weight of mixed, masked, and contrastive objectives, respectively. This hybrid strategy strengthens token representations for downstream tasks.
4.4 Efficient Fine-tuning
Fine-tuning LLMs requires balancing computational efficiency with task-specific performance. This challenge is best addressed through a co-design approach, integrating model-level techniques (e.g., parameter-efficient updates, sparsity) with system-level optimizations (e.g., memory compression, resource allocation). By jointly optimizing these layers, fine-tuning methods can achieve scalability and performance under constrained resources.
4.4.1 Efficient Fine-Tuning for Parameters
Parameter-efficient fine-tuning (PEFT) techniques, such as Low-Rank Adaptation (LoRA) and Prefix-Tuning, aim to reduce the number of trainable parameters while maintaining model performance. These methods modify the training process to focus on smaller, task-specific updates instead of adjusting all model parameters (Liu et al., 2024d; Zhang et al., 2023b; Liu et al., 2021a; He et al., 2023; Zhang et al., 2023a; Ding et al., 2023).
LoRA reduces memory and computation requirements by introducing a low-rank decomposition for weight updates (Liu et al., 2024d; Zhang et al., 2023b),. The model weights $W\in\mathbb{R}^{d\times d}$ are kept frozen, while updates are parameterized as $\Delta W=\mathbf{A}\mathbf{B}^{\top}$, where $\mathbf{A},\mathbf{B}\in\mathbb{R}^{d\times r}$ and $r\ll d$. Training only $\mathbf{A}$ and $\mathbf{B}$ minimizes resource usage while preserving the model’s capacity to adapt to new tasks. To further optimize resource utilization, system-level techniques such as quantization are applied to $\mathbf{A}$ and $\mathbf{B}$. Quantization scales the matrices using step sizes (e.g., $\Delta_{\mathbf{A}}$ and $\Delta_{\mathbf{B}}$), enabling hardware-friendly computation with reduced memory overhead (Lv et al., ).
Prefix-Tuning adapts the model by learning a continuous prefix embedding $P\in\mathbb{R}^{k\times d}$, which is prepended to the input embeddings. The pre-trained model weights remain unchanged. The optimization objective for Prefix-Tuning is:
$\min_{P}\mathcal{L}_{\text{task}}(f_{\theta}([P;X]),Y),$
where $X$ is the input, $Y$ is the target, and $f_{\theta}$ is the frozen pre-trained model. This method allows efficient adaptation to a wide range of tasks with minimal computational costs (Li and Liang, 2021; Zhang et al., 2023a; Ding et al., 2023).
System-level optimizations can further enhance the efficiency of Prefix-Tuning. For example, the prefix matrix $P$ can be compressed using a low-rank approximation, where $P\approx\mathbf{U}_{P}\mathbf{V}_{P}^{\top}$ and $\mathbf{U}_{P},\mathbf{V}_{P}\in\mathbb{R}^{k\times r}$. This reduces memory usage by storing only the decomposed components. The optimization objective then becomes:
$\min_{\mathbf{U}_{P},\mathbf{V}_{P}}\|\mathcal{F}_{\theta}([P;X])-\mathcal{F}_{\theta}([\mathbf{U}_{P}\mathbf{V}_{P}^{\top};X])\|_{F}^{2}+\lambda\cdot\mathcal{C}(\mathbf{U}_{P},\mathbf{V}_{P}),$
where $\mathcal{C}(\mathbf{U}_{P},\mathbf{V}_{P})$ represents the memory cost of the decomposed components. This coupling of model compression and system-aware design ensures that Prefix-Tuning remains efficient across diverse hardware platforms (Liu et al., 2021a; Lv et al., ).
By integrating low-rank decompositions and system-level optimizations, PEFT techniques demonstrate the principles of model-system co-design. These methods enable scalable fine-tuning for large models while ensuring efficient resource utilization. LoRA and Prefix-Tuning have proven effective in adapting large language models to specific tasks with minimal memory and computational costs (Liu et al., 2024d; Zhang et al., 2023b; Li and Liang, 2021; Liu et al., 2021a).
4.4.2 Efficient Fine-Tuning for Sparsity
Sparse fine-tuning leverages inherent sparsity in gradients and parameter updates to optimize computational and memory efficiency, particularly in the context of large-scale language models. Unlike traditional fine-tuning approaches that involve dense updates across all model parameters, sparse fine-tuning identifies and updates only the most impactful weights or gradients, significantly reducing resource consumption while preserving task performance.
Scaling Sparse Fine-Tuning (SSF) dynamically applies sparsity to the gradient matrix $G\in\mathbb{R}^{d\times d}$ by introducing a binary mask $M\in\{0,1\}^{d\times d}$, which selects critical elements based on a combination of gradient magnitude and computational cost. Specifically, the mask $M$ is defined such that $M_{ij}=1$ if $|G_{ij}|/\text{Cost}(G_{ij})\geq\tau$, where $\tau$ is an adaptive threshold adjusted during training. The resulting sparse update is then computed as $\Delta W=M\odot G$, where $\odot$ denotes element-wise multiplication (Ansell et al., 2024). By incorporating cost metrics like memory bandwidth or latency, this method aligns gradient sparsity with hardware constraints.
Layerwise Importance Sampling for Memory-Efficient Fine-Tuning (LISA) further optimizes sparse fine-tuning by focusing on the most critical layers of the model. Each layer’s importance is evaluated using the gradient norm or its contribution to the task-specific loss, $\mathcal{L}_{\text{task}}$. Formally, the layer subset $\mathcal{S}$ to be updated is selected as:
$\mathcal{S}=\operatorname*{arg\,max}_{\mathcal{S}\subseteq\{1,\ldots,L\}}\sum_{l\in\mathcal{S}}\frac{\|\nabla_{\theta^{(l)}}\mathcal{L}_{\text{task}}\|_{F}}{\text{Cost}(\nabla_{\theta^{(l)}})},\quad|\mathcal{S}|\leq k,$
where $\nabla_{\theta^{(l)}}\mathcal{L}_{\text{task}}$ denotes the gradient of the loss with respect to layer $l$’s parameters, and $\text{Cost}(\cdot)$ represents the resource overhead of updating that layer. This approach ensures that updates are concentrated on the most impactful layers while adhering to predefined resource budgets (Pan et al., 2024).
In addition to layerwise optimization, sparse fine-tuning benefits from system-level enhancements like gradient caching and asynchronous execution. For example, momentum-based gradient sparsity propagates prior updates to stabilize training under sparsity constraints. Let $\Delta W^{(t-1)}$ be the previous sparse update. At step $t$, the new sparse update integrates past momentum as:
$\Delta W^{(t)}=M^{(t)}\odot\left(G^{(t)}+\beta\cdot\Delta W^{(t-1)}\right),$
where $\beta$ is a momentum hyperparameter that smooths the gradient trajectory across sparse updates. This formulation ensures consistent parameter updates despite high sparsity (Dettmers and Zettlemoyer, 2019; Waleffe et al., 2024).
Furthermore, system optimizations such as asynchronous execution enhance scalability in distributed environments. Sparse gradient updates are computed independently across $N_{\text{devices}}$ processing units, with synchronization occurring intermittently. If $T_{\text{comp}}$ is the computation time for sparse updates and $T_{\text{sync}}$ the synchronization overhead, the total time per step is: $T_{\text{total}}=\max\left(T_{\text{comp}},\frac{T_{\text{sync}}}{N_{\text{devices}}}\right)$, where overlapping computation and communication minimizes total latency (Sheng et al., 2023; Huang et al., 2019).
Sparse fine-tuning techniques have demonstrated broad applicability, ranging from domain-specific tasks (e.g., legal and biomedical text generation) to resource-constrained environments, such as edge devices or distributed training. Future directions may explore tighter integration between sparse fine-tuning and hardware-level optimizations, including tensor sparsity on GPUs and custom accelerators tailored for sparse computation (Ansell et al., 2024; Pan et al., 2024; Liu et al., 2024d; Zhang et al., 2023b).
4.5 Model-System Co-Design Unified Framework
Efficient pretraining leverages model and system optimizations to balance accuracy, memory, and computational efficiency. A unified framework integrates key aspects of quantization, sparsity, and distributed system design to maximize efficiency. The objective is to optimize pretraining by explicitly modeling system constraints and pretraining-specific features.
Let $\mathcal{L}_{\text{task}}(\theta)$ represent the pretraining loss, where $\theta$ includes parameters such as weights $W$, quantization factors $\Delta$, and sparsity masks $\mathbf{M}$. To enhance pretraining efficiency, weights are quantized and sparse updates are applied. Quantized weights are expressed as $W=\mathcal{Q}(\hat{W},\Delta)$, where $\mathcal{Q}(\hat{W},\Delta)$ denotes quantization with step size $\Delta$. Sparse masks $\mathbf{M}$ restrict updates to the most critical elements, ensuring computational efficiency.
The unified objective integrates the pretraining loss with quantization and sparsity constraints:
$\mathcal{L}_{\text{pretrain}}=\sum_{i=1}^{L}\|\mathbf{M}_{i}\odot\mathcal{Q}(\hat{W}_{i},\Delta_{i})\|_{F}^{2}+\sum_{j=1}^{n}\|\mathbf{M}_{A,j}\odot\mathcal{F}_{\theta}(X_{j})-Y_{j}\|^{2},$
where: - $\mathbf{M}_{i}$ represents the weight sparsity mask for layer $i$. - $\mathbf{M}_{A,j}$ denotes the attention sparsity mask for token $j$. - $\mathcal{F}_{\theta}(X_{j})$ is the model output for input $X_{j}$, compared to target $Y_{j}$.
5 Conclusion
Efficient Foundation Model Design techniques aim to build models that are computationally efficient while achieving faster training and inference speeds without sacrificing performance. These techniques involve three main areas: model design, system design, and model-system co-design.
Model design focuses on optimizing the internal structure of models for acceleration. Quantization provides a hardware-focused perspective, compressing models into lower-precision representations to optimize storage and computation. Knowledge distillation allows efficient knowledge transfer, enabling smaller student models to study the performance from larger teacher models. Pruning offers a neural network model reduction by eliminating redundant connections.
System design introduces optimizations at the system design infrastructure level, it include KV cache compression, parallelism, contextual sparsity, etc. KV cache compression minimizes memory usage by reducing the storage of key-value caches, which is efficient in long-context processing. Parallelism improves throughput by distributing computation across multiple GPUs/CPUs. Sparsity eliminates unnecessary computations in sparse attention matrix, which further reduces overhead. Low-level GPU optimizations maximize hardware utilization, which can achieve faster and more scalable model execution.
Model-system co-design bridges the gap between model and system design. MoE enables dynamic routing to reduce computation while maintaining performance for diverse tasks. Fine-tuning techniques enhance adaptability on downstream tasks. Mixed-precision training reduces computation by leveraging both low- and high-precision computation. These model and system co-design strategies leads to better resource utilization and faster inference.